3. Four students are to be selected for the leadership group of a club that currently has 20 students. Of the 20 students in the club, there are 9 CS majors, 5 English majors, 3 Sociology majors, 2 Math majors, and 1 Data Science major (a) How many different groups of 4 students are possible, assuming order does not matter (and majors do not matter)? (b) If each group is equally likely to be picked, what is the probability that the Data Science major will be in the selected group? (c) What is the probability that at least one English Major will be in the group? (d) If order did matter (say because the first person picked will be the President of the club, the 2nd person will be the VP, et cetera), how many different groups of four students can be made??
Answer #1

Solution

Back-up Theory

Number of ways of arranging n things among themselves = n! = n(n - 1)(n - 2) …… 3.2.1.................................….(1)

Values of n!can be directly obtained using Excel Function: Math & Trig FACT (Number)...................................... (1a)

Number of ways of selecting r things out of n things is given by nCr = (n!)/{(r!)(n - r)!}......................................……(2)

Values of nCr can be directly obtained using Excel Function: Math & Trig COMBIN(Number, Number_chosen)

[Number is n, Number_chosen is r]………………….................................................................……………………. (2a)

Probability of an event E, denoted by P(E) = n/N ………………………………………….......................................…(3)

where n = n(E) = Number of outcomes/cases/possibilities favourable to the event E and N = n(S) = Total number all possible outcomes/cases/possibilities.

Now to work out the solution,

Part (a)

Given that order does not matter and majors also do not matter, the situation is one of selecting any 4 students out of 20 students. This, vide (2), is 20C4 = (20!)/{(4!)(16)!}

= (20 x 19 x 18 x 17)/(4 x 3 x 2 x 1)

= 4845 Answer

Part (b)

Since there is only one Data Science Major student in the club, the inclusion of that student can be done only in one way. Once this is ensured, we need to select three from the remaining 19, which, vide (2), can be done 19C3 = 969 ways [vide (2a)].

Thus, vide (3), n = 1 x 969 = 969.

Vide (3) again, and answer of Part (a), N = 4845. So, the required probability

= 969/4845

= 0.2   Answer

Part (c)

Probability at least one English Major is in the group

= 1 – P(Group does not include any English Major student)

= 1 – P(All 4 members are selected out of the remaining 15, i.e., excluding 5 English Major students)

= 1 – {(15C4)/( 20C4)} [identical to the above two analyses]

= 1 – {1365/4845} [vide (2a)]

= 1 – 0.2817

= 0.7183 Answer

Part (d)

If the order matters, this becomes a permutation problem of arranging 20 taking 4 at a time, which can be done in (20 x 19 x 18 x 17) ways

= 116280 Answer

DONE

[Going beyond, answer to Part (d) is also equal to {answer to Part (a) x 4!}]